OpenStudy (anonymous):

can some1 tell me if im right with this problem

OpenStudy (anonymous):

\[\frac{ 5 }{ a }-\frac{ 3 }{ b }+\frac{ 1 }{ ab }\] that's the problem I got \[\frac{ 5b-3a+1 }{ ab }\]